OpenStudy (anonymous):

The length of a rectangle is 3cm longer than its width. Its wiidth is x cm. Its perimeter is 54 cm. Write a equtation that represents the sentence. Answers are: a 3x=54 b.2x+2(x+3)=54 c.2x+3=54 d.x+3=54 i think it is a, am i right?

OpenStudy (anonymous):

No, the answer is B because perimeter is twice the width added to twice the length and twice x and twice (x+3)=54

OpenStudy (anonymous):

width is x so length is x+ 3 perimeter = 2x + 2(x+ 3) = 54 b